OpenStudy (anonymous):

Write the slope-intercept form of the equation parallel to y = 7x + 2, which passes through the point (1, -3). A. y = 7x - 10 B. y = -1/7 x + 2 C. y = 7x - 3 D. y = -7x + 10 Write the point-slope form of the equation passing through (5, -1) with a slope of 6. A. y-1=6(x+5) B. y+5=6(x-1) C. y+1=6(x-5) D. y-5=6(x+1) Write the equation for the vertical line passing through (5, 8). A. x = 5 B. x = 8 C. x = 0 D. x = 5/8 Write the equation for the horizontal line passing through (4, -6). A. y = 4 B. y = -6 C. y = 4/6 D. y = -6/4

OpenStudy (anonymous):

it's multiple choice take what are your answers?

OpenStudy (anonymous):

1. parallel lines have the same slope (x). 2. y - y1 = m(x - x1) is point slope form, to use it simply plug in the x and y values of a coordinate into x1 and y1

OpenStudy (anonymous):

3. (5, 8) what is the x value? 4. (4, -6) what is the y value? atleast show some effort :P

OpenStudy (anonymous):

im terrible at math

OpenStudy (anonymous):

practice and you'll get better. us showing you the answer will only set you up for failure when the test comes around.

OpenStudy (anonymous):

what do you have trouble with?

OpenStudy (anonymous):

this kind of stuff

OpenStudy (anonymous):

ok, let's start with #1 Write the slope-intercept form of the equation parallel to y = 7x + 2, which passes through the point (1, -3).

OpenStudy (anonymous):

first: what is slope intercept form?

OpenStudy (anonymous):

y=mx+b?

OpenStudy (anonymous):

yep and in y = mx + b. m = slope and b is a number.

OpenStudy (anonymous):

lines are parallel if they have the same slope. in english: if the value for m is equal in two equations, then they have the same slope and are parallel.

OpenStudy (anonymous):

one more thing to address: have you heard of point slope form?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

great, in a problem like this you'll need to use point slope form which is: y - y1 = m(x - x1) x1 = x coordinate y1 = y coordinate m = slope

OpenStudy (anonymous):

go back to what you were given: y = 7x + 2 and coordinates (1, -3). plug them into point slope form. in the equation "y = 7x + 2" you only need the slope.

OpenStudy (anonymous):

still confused? i'll write it out

OpenStudy (anonymous):

ya just a little

OpenStudy (anonymous):